OpenStudy (anonymous):

what is the length of the major axis of the conic section below? ((x-3)^2/(81))+((y+6)^2/(49))=1

OpenStudy (aum):

Compare the given equation to the standard form:\[ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1 \]The length of the major axis is 2a. So compare the two equations, find 'a', and double it to find the length of the major axis.

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

a^2 in your equation is 81, take the square root of that and you get 9. Length of the major axis is 2*a. Which is 2*9=18.

OpenStudy (anonymous):

that make a lot of since. thank you!!!

OpenStudy (anonymous):

i don't understand